### v1

@@ 6 The classical Poincaré series
Any collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=M_{\varkappa}$ obviously belongs to $A_{j+1}$. According to Lemma 5, $a_{M_{\varkappa}}^{(j+1)}\leq[\mathbb{K}_{j}:\mathbb{K}_{0}]$. Let us take the collection $\{k^{0}_{\sigma}\}$ with $k^{0}_{\varkappa}=1$ and $k^{0}_{\sigma}=0$ for $\sigma\neq\varkappa$. One has $\nu(\{k^{0}_{\sigma}\})=M_{\varkappa}$. Let $f=0$ be an equation of a $G_{\varkappa}(=G_{j})$-curvette intersecting the component $E_{\varkappa}$ at the point with the coordinate $\alpha\in\mathbb{K}_{j}$. Here we assume that the coordinate is equal to zero at the intersection point with the component on the geodesic to $\delta_{C}$ and to $\infty$ at the one with the component on the geodesic to $\sigma_{0}$. Let $g$ be one of these functions, say, with $\alpha=1$. The ratio $\Psi={\widetilde{f}}/{\widetilde{g}}$, for $f\in{\mathcal{E}}^{\{k^{0}_{\sigma}\}}_{\mathbb{K}^{2},0}$, has no zeroes or poles on all the components of each of the indicated geodesics and thus is constant on each of them. On the component $E_{\varkappa}$ this ratio is of the form $c\frac{w-\alpha}{w-1}$ with $c\in\mathbb{K}$ and (arbitrary!) $\alpha\in\mathbb{K}_{j}$. The value on $\delta_{C}$ (equal to the value at $w=0$, i.e. $c\alpha$) is an arbitrary point of $\mathbb{K}_{j}$. Therefore $d^{\{k^{0}_{\sigma}\}}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$ and $a_{M_{\varkappa}}^{(j+1)}\geq[\mathbb{K}_{j}:\mathbb{K}_{0}]$. ∎
@@
If $v\in S^{\mathbb{K}}_{C}$ cannot be represented as $M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, one has $a_{v}^{j+1}=a_{v}^{j}$, what is compatible with (10). Let $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, $11\leq\ell<\ell_{j}-$ 1, and $v$ cannot be represented as $(\ell+1)M_{\rho_{j}}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$. In this case Equation (10) reduces to
@@
(all other summands are equal to zero). Since, for $10\leq\ell^{\prime}\leq\ell-$ 1, $v-\ell^{\prime}M_{\rho_{j}}=M_{\rho_{j}}+b^{*}$ with $b^{*}\in S^{\mathbb{K}}_{C}$ and, due to Lemma 7, $a_{M_{\rho_{j}}}^{(j)}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$, all the summands in (11) except the last one are equal to $[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$. Thus in this case (10) follows from Lemma 6.
@@
    The value $v$ can be represented as $M_{\varkappa}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$, $\varkappa>\rho_{j}$. In this case (10) follows from Lemmas 7 and  8.
@@
Propositions 2 and 3 imply the statement of Theorem 2 in the case under consideration: $\rho_{1}\neq\sigma_{0}$. In the case $\rho_{1}=\sigma_{0}$, Lemmas 58 and thus Proposition 3 hold for $j\geq 2$. Proposition 2 does not hold. However, it is possible to compute the series $P^{(2)}_{C}(t)$ directly using Lemmas 58. Thus the statement of Theorem 2 holds in this case as well. ∎
@@
Any series $\zeta(t)=1+c_{1}t+c_{2}t^{2}+\cdots\in\mathbb{Z}[[t]]$ can be in a unique way represented as $\zeta(t)=\prod_{m=1}^{\infty}(1-t^{m})^{s_{m}}$ with $s_{m}\in\mathbb{Z}$. We say that the series is cyclotomic if this product/ratio of binomials is finite. There is a natural generalization of this notion for series in sevearl variables. Ussually in previous computations (say e.g. [5]) the Poincaré series appeared to be cyclotomic. This is not the case for the classical Poincaré series $P_{C}(t)$ in case II, i.e. when $s=\infty$.
@@
Let us assume that the curve valuation is of type I. Let $c$ be the conductor of the semigroup $S^{\mathbb{K}}_{C}$ (i.e the minimal element of $S^{\mathbb{K}}_{C}$ such that $c+\mathbb{Z}_{\geq 0}\subset S^{\mathbb{K}}_{C}$). It is known that $c$ can be expresed as
@@
Let $\Delta:=c+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}$ and $\ell=\ell_{1}\cdots\ell_{s}=[\mathbb{K}_{s}:\mathbb{K}]$. The computation of the Poincaré series gives that $a_{v}=\dim J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)=\ell$ for all $v\geq\Delta$ and $a_{\Delta-1}<\ell$. Moreover, one has the following symmetry property for all $v\in\mathbb{Z}$:
$\dim\left({J^{\mathbb{K}}(v)}\left/{J^{\mathbb{K}}(v+1)}\right.\right)+\dim\left({J^{\mathbb{K}}(\Delta-1-v)}\left/{J^{\mathbb{K}}(\Delta-v)}\right.\right)=\ell\;$ (13)
Following the notations of Remark 2, the integral closure $\overline{R}$ of the local ring $R=\mathbb{K}[[x,y]]/(F)$ of the algebroid curve defined by $F$ is a valuation ring and is isomophic to $\mathbb{K}^{\prime}[[u]]$ ($u$ is an uniformizing parameter). One has:
@@
Let
$\Delta=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}\;.$
Then, the conductor ideal ${\mathcal{C}}(\overline{R}/R)=\{z\in\overline{R}:z\overline{R}\subset R\}$ is the ideal $u^{\Delta}\overline{R}$.
Note that the ring $R$ is Gorenstein (being plane) and the symmetry of the equation (13) is the just the characterization of the Gorenstein condition of the ring $R$ (cf. [3]).
@@ 7 The Poincaré series of a divisorial valuation
Let $\nu=\nu_{\delta}$ be a divisorial valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$. This means that it is defined by a component $E_{\delta}$ of the exceptional divisor of a modification of $(\mathbb{C}^{2},o)$. Let $\pi:(W,{\mathcal{D}})\to(\mathbb{C}^{2},0)$ be the minimal $\mathbb{K}$-resolution of the valuation $\nu$. The quotient of the dual graph of this resolution by the action of the Galois group $G$ looks like in figure 6.
@@
This graph essentially coincides with the minimal resolution graph of a $\mathbb{K}_{\delta}$-curvette at the component $E_{\delta}$ ($\mathbb{K}_{\delta}=\mathbb{K}_{s}$) with a possible tail from $\tau_{g}$ to $\delta$ added. Here, as above, $\sigma_{i}$, $i=0,1,\ldots,g$ are the dead ends of the graph $\check{\Gamma}$, $\tau_{i}$, $i=1,\ldots,g$ are the rupture points and $\rho_{j}$, $j=1,\ldots,s$ are the branching points of the $\mathbb{K}$-resolution.
@@
We shall keep the notations ${\mathcal{E}}^{\{k_{\sigma}\}}_{\mathbb{K},0}$, $F^{\{k_{\sigma}\}}$, and $d^{\{k_{\sigma}\}}$ as before.
@@
$P_{\nu}(t)=P^{S}(t)\frac{1}{1-t^{M_{\delta}}}\cdot\prod_{i=1}^{s}\frac{1-t^{\ell_{i}M_{\rho_{i}}}}{1-t^{M_{\rho_{i}}}}\;.$ (14)
@@
Let $C$ be a $\mathbb{K}_{s}$-curvette at the component $E_{\delta}$. For the coefficient $a_{v}$ in the Poincaré series $P_{\nu}(t)=\sum a_{v}t^{v}$ one has $a_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximun is taken over the collections $\{k_{\sigma}\}$ such that $\nu(\{k_{\sigma}\})=\sum k_{\sigma}M_{\sigma}=v$. Let $A^{\prime}$ be the set of collections $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$ such that $k_{\delta}=0$ and let $P^{\prime}(t)=\sum a^{\prime}_{v}t^{v}$ with $a^{\prime}_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximun is taken over the collections $\{k_{\sigma}\}\in A^{\prime}$ such that $\nu(\{k_{\sigma}\})=v$.
@@
$P^{\prime}_{\nu}(t)=P_{C}(t)=P^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}\;.$
@@
Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\delta}+b$ with $b\in S^{\mathbb{K}}_{C}$ and the maximal possible $\ell$ in this representation. Then $a_{v}=\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$.
@@
(Cf. proof of Lemma 6.) Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $k_{\sigma}=k^{1}_{\sigma}+k^{2}_{\sigma}$, where $k^{1}_{\delta}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\delta$, $k^{2}_{\delta}=0$. A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$. Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $h=h_{1}h_{2}$. As in Lemma 6, we shall use a coordinate on $E_{\delta}$ such that the intersection point with the previous component from the geodesic to $\sigma_{0}$ is equal to $\infty$ and we shall take the function $h_{1}$ described there (that is, the $\ell^{\prime}$th power of the left hand side of the equation of the $G_{s}$-curvette at the point $w=1$ of the component $E_{\delta}$). One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$. The function $\Psi_{2}$ is constant on the component $E_{\delta}$. The maximal possible dimension (over $\mathbb{K}$) of the set of its values on this component is $a^{\prime}_{b^{\prime}}$ ($b^{\prime}=v-\ell^{\prime}M_{\delta}$). The function $\Psi_{1}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesic between $\sigma_{0}$ and $\delta$, excluding the last one. Its restriction to the component $E_{\delta}$ is equal to $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})/(w-1)^{\ell^{\prime}}$ with $c\in\mathbb{K}$, where $w_{p}\in\mathbb{K}_{s}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\delta}$. Therefore, up to the fixed multiple $(w-1)^{\ell^{\prime}}$, the restriction of $\Psi$ to the component $E_{\delta}$ is equal to the restriction of $\Psi_{2}$ multiplied by $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})$. The dimension (over $\mathbb{K}$) of the space of functions of the latter form is $\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}]$ (since $w_{p}$ are arbitrary elements of $\mathbb{K}_{s}$). Therefore the (maximal) dimension of the set of these functions on $E_{\delta}$ is $a^{\prime}_{b^{\prime}}+\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}_{0}]$. Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$. ∎
@@
$P_{\nu}(t)=P^{\prime}_{\nu}(t)\cdot\frac{1}{1-t^{M_{\delta}}}.$
@@
One can see that the classical Poincaré series of a curve or of a divisorial valuation determines the combinatorial type of the minimal $\mathbb{K}$-resolution of the valuation.

### v2

@@ 6 The classical Poincaré series of a curve valuation
Any collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=M_{\varkappa}$ obviously belongs to $A_{j+1}$. According to Lemma 4, $a_{M_{\varkappa}}^{(j+1)}\leq[\mathbb{K}_{j}:\mathbb{K}_{0}]$. Let us take the collection $\{k^{0}_{\sigma}\}$ with $k^{0}_{\varkappa}=1$ and $k^{0}_{\sigma}=0$ for $\sigma\neq\varkappa$. One has $\nu(\{k^{0}_{\sigma}\})=M_{\varkappa}$. Let $f=0$ be an equation of a $G_{\varkappa}(=G_{j})$-curvette intersecting the component $E_{\varkappa}$ at the point with the coordinate $\alpha\in\mathbb{K}_{j}$. Here we assume that the coordinate is equal to zero at the intersection point with the component on the geodesic to $\delta_{C}$ and to $\infty$ at the one with the component on the geodesic to $\sigma_{0}$. Let $g$ be one of these functions, say, with $\alpha=1$. The ratio $\Psi={\widetilde{f}}/{\widetilde{g}}$, for $f\in{\mathcal{E}}^{\{k^{0}_{\sigma}\}}_{\mathbb{K}^{2},0}$, has no zeroes or poles on all the components of each of the indicated geodesics and thus is constant on each of them. On the component $E_{\varkappa}$ this ratio is of the form $c\frac{w-\alpha}{w-1}$ with $c\in\mathbb{K}$ and (arbitrary!) $\alpha\in\mathbb{K}_{j}$. The value on $\delta_{C}$ (equal to the value at $w=0$, i. e., $c\alpha$) is an arbitrary point of $\mathbb{K}_{j}$. Therefore $d^{\{k^{0}_{\sigma}\}}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$ and $a_{M_{\varkappa}}^{(j+1)}\geq[\mathbb{K}_{j}:\mathbb{K}_{0}]$. ∎
@@
If $v\in S^{\mathbb{K}}_{C}$ cannot be represented as $M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, one has $a_{v}^{j+1}=a_{v}^{j}$, what is compatible with (14). Let $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, $1\leq\ell<\ell_{j}-1$, and $v$ cannot be represented as $(\ell+1)M_{\rho_{j}}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$. In this case Equation (14) reduces to
@@
(all other summands are equal to zero). Since, for $0\leq\ell^{\prime}\leq\ell-1$, $v-\ell^{\prime}M_{\rho_{j}}=M_{\rho_{j}}+b^{*}$ with $b^{*}\in S^{\mathbb{K}}_{C}$ and, due to Lemma 6 (Lemma 3 for $j=1$), $a_{M_{\rho_{j}}}^{(j)}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$, all the summands in (15) except the last one are equal to $\mathbb{K}_{j-1}:\mathbb{K}_{0}]$. Thus in this case (14) follows from Lemma 5.
@@
    The value $v$ can be represented as $M_{\varkappa}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$, $\varkappa>\rho_{j}$. In this case (14) follows from Lemmas 6 and  7.
@@
Proof of Theorem 2. If $\rho_{1}\neq\sigma_{0}$, the statement follows directly from Propositions 3 and 4. If $\rho_{1}=\sigma_{0}$, Lemmas 47 and thus Proposition 4 hold for $j\geq 2$. Proposition 3 does not hold. However, it is possible to compute the series $P^{(2)}_{C}(t)$ directly using Lemmas 47. Thus the statement of Theorem 2 holds in this case as well. $\Box$
@@
Any series $\zeta(t)=1+c_{1}t+c_{2}t^{2}+\cdots\in\mathbb{Z}[[t]]$ can be in a unique way represented as $\zeta(t)=\prod_{m=1}^{\infty}(1-t^{m})^{s_{m}}$ with $s_{m}\in\mathbb{Z}$. We say that the series is cyclotomic if this product/ratio of the binomials is finite. There is a natural generalization of this notion for series in several variables. Usually in previous computations (say, e.g., [4]) the Poincaré series appeared to be cyclotomic. This is not so for the Poincaré series $P^{\mathbb{K}}_{C}(t)$ in case II, i. e., when $s=\infty$.
@@
Let us assume that the curve valuation corresponds to Case I. Let $c$ be the conductor of the semigroup $S^{\mathbb{K}}_{C}$ (i.e., the minimal element of $S^{\mathbb{K}}_{C}$ such that $c+\mathbb{Z}_{\geq 0}\subset S^{\mathbb{K}}_{C}$). It is known that $c$ can be expressed as
@@
Let $\ell=\ell_{1}\cdots\ell_{s}=[\mathbb{K}_{s}:\mathbb{K}]$ and
$\Delta:=c+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}\;.$
The computation of the Poincaré series gives that $a_{v}=\dim J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)=\ell$ for all $v\geq\Delta$ and $a_{\Delta-1}<\ell$. Moreover, one has the following symmetry property:
@@
For all $v\in\mathbb{Z}$ one has
[…]
Following the notations of Remark 4, the integral closure $\overline{R}$ of the local ring $R=\mathbb{K}[[x,y]]/(F)$ of the algebroid curve defined by $F$ is a valuation ring and is isomorphic to $\mathbb{K}^{\prime}[[u]]$ ($u$ is an uniformizing parameter). One has:
Corollary 2.
The conductor ideal ${\mathcal{C}}(\overline{R}/R)=\{z\in\overline{R}:z\overline{R}\subset R\}$ is the ideal $u^{\Delta}\overline{R}$.
Note that the ring $R$ is Gorenstein (being plane) and the symmetry described by the equation (17) is just the characterization of the Gorenstein condition of the ring $R$ (cf. [3]).
@@ 7 The Poincaré series of a divisorial valuation
Let $\nu=\nu_{\delta}$ be a divisorial valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$. This means that it is defined by a component $E_{\delta}$ of the exceptional divisor of a modification of $(\mathbb{C}^{2},0)$. We assume that the corresponding $G$-invariant modification is finite. Otherwise we are in a situation similar to Case III in the curve case and the valuation $\nu_{\delta}$ is a multiple of another one: corresponding to the component on which the $G$-orbit of the blown-up point is infinite. Let $\pi:(W,{\mathcal{D}})\to(\mathbb{C}^{2},0)$ be the minimal $G$-resolution of the valuation $\nu$. The quotient of the dual graph of this resolution by the action of the Galois group $G$ looks like in figure 7.
@@
This graph essentially coincides with the minimal resolution graph of a $\mathbb{K}_{\delta}$-curvette at the component $E_{\delta}$ ($\mathbb{K}_{\delta}=\mathbb{K}_{s}$) with a possible tail from $\tau_{g}$ to $\delta$ added. Here, as above, $\sigma_{i}$, $i=0,1,\ldots,g$, are the dead ends of the graph $\check{\Gamma}$, $\tau_{i}$, $i=1,\ldots,g$, are the rupture points and $\rho_{j}$, $j=1,\ldots,s$, are the splitting points of the $G$-resolution.
@@
We shall keep the notations ${\mathcal{E}}^{\{k_{\sigma}\}}_{\mathbb{K},0}$, $F^{\{k_{\sigma}\}}$, and $d^{\{k_{\sigma}\}}$ as before. Let $C$ be a $\mathbb{K}_{s}$-curvette at the component $E_{\delta}$.
@@
$P^{\mathbb{K}}_{\nu_{\delta}}(t)=P^{S}_{C}(t)\cdot\frac{1}{1-t^{M_{\delta}}}\cdot\prod_{i=1}^{s}\frac{1-t^{\ell_{i}M_{\rho_{i}}}}{1-t^{M_{\rho_{i}}}}\;.$ (18)
@@
For the coefficient $a_{v}$ in the Poincaré series $P^{\mathbb{K}}_{\nu_{\delta}}(t)=\sum a_{v}t^{v}$ one has $a_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximum is taken over the collections $\{k_{\sigma}\}$ such that $\nu(\{k_{\sigma}\})=\sum k_{\sigma}M_{\sigma}=v$. Let $A^{\prime}$ be the set of collections $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$ such that $k_{\delta}=0$ and let $P^{\prime}(t)=\sum a^{\prime}_{v}t^{v}$ with $a^{\prime}_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximum is taken over the collections $\{k_{\sigma}\}\in A^{\prime}$ such that $\nu(\{k_{\sigma}\})=v$.
@@
$P^{\prime}_{\nu}(t)=P_{C}^{\mathbb{K}}(t)=P^{S}_{C}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}\;.$
@@
Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\delta}+b$ with $b\in S^{\mathbb{K}}_{C}$ and such that $b-M_{\delta}\notin S_{C}^{\mathbb{K}}$. Then $a_{v}=\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$.
@@
[…]
$\frac{c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})}{(w-1)^{\ell^{\prime}}}$
with $c\in\mathbb{K}$, where $w_{p}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\delta}$. The set of these points is an arbitrary union of $G_{s}$-orbits. Therefore, up to the fixed multiple $(w-1)^{\ell^{\prime}}$, the restriction of $\Psi$ to the component $E_{\delta}$ is equal to the restriction of $\Psi_{2}$ multiplied by $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})$. The dimension (over $\mathbb{K}$) of the space of functions of the latter form is $\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}]$ (since the collection of the symmetric polynomials of the points $w_{p}$ is a generic point of $\mathbb{K}_{s}^{\ell^{\prime}}$). Therefore the (maximal) dimension of the set of these functions on $E_{\delta}$ is $a^{\prime}_{b^{\prime}}+\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}_{0}]$. Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$. ∎
@@
$P^{\mathbb{K}}_{\nu_{\delta}}(t)=P^{\prime}(t)\cdot\frac{1}{1-t^{M_{\delta}}}=P^{\mathbb{K}}_{C}(t)\cdot\frac{1}{1-t^{M_{\delta}}}.$
@@
One can see that the Poincaré series of a curve or of a divisorial valuation determines the combinatorial type of the minimal $G$-resolution of the valuation.
Remark 15.
The formula in Proposition 5 relating the Poincaré series of the divisorial valuation $\nu$ and the Poincaré series of a $\mathbb{K}_{s}$-curvette $C$ at the component $E_{\delta}$ obviously extends the known one for the complex case (see [12] or [10]). Even the formula itself and, to some extent, the proofs have some technical resemblance with the relation of both series in the case of several divisorial valuations in [9]. However, the objects are rather diferent, in [9] all the considerations are in the pure complex analytic framework and the proofs use only properties of pencils (another way to treat meromorphic functions like in this paper) in the complex setting.